what is the the slope intercept form for x-5y=0
Answer by peter_mccarter(5)

Y= mx+b. This is the slope intercept form. What is needed here is to rearrange the equation to the form above so x-5y=0, we move the x to the right and have -5y=-x+0. Now it is in the slope intercept form. To solve you must do more steps multiply by -1 and have 5y=x+0 now divide the equation by 5 to get the slope y= mx=b /b is the y intercept/ /m is the slope/ the final answer is y=1/5x+0 slope is 1/5 and intercept is (0,0).